OpenStudy (geekfromthefutur):

Barry has been watching the geese that live in his neighborhood. The number of geese changes each week. n f(n) 1 56 2 28 3 14 4 7 Which function best shows the relationship between n and f(n)? f(n) = 28(0.5)n f(n) = 56(0.5)n-1 f(n) = 56(0.5)n f(n) = 112(0.5)n-1

OpenStudy (jack1):

the easiest way to check is with n=1 first, then confirm that ur answer holds true for the other values of n so first try: \( \large f(n) = 28(0.5)^n\) for n=1 \( \large f(n) = 28(0.5)^1\) \( \large f(n) = 28 \times 0.5 = 14 ....\) f(n) for n = 1 is supposed to equal 56... so A is not ur answer... now u try with B

OpenStudy (shamim):

just put the value of n and put the value of f(n) in each equation.

OpenStudy (jack1):

so @geekfromthefutur how did u go with option B... did it work or did u move on to C...?

OpenStudy (geekfromthefutur):

it had worked

OpenStudy (jack1):

sweet, then B is ur answer... but still confirm with n=2 to make sure, yeah?

OpenStudy (geekfromthefutur):

yea :3 thx @Jack1

OpenStudy (jack1):

no worries ;)